OpenStudy (anonymous):

What is the slope of a line that passes through the point (−2, 3) and is parallel to a line that passes through (3, 7) and (−2, −8)?

OpenStudy (3psilon):

Use the slope formula with the two last points \[\frac{ y_{2 } - y_{1}}{ x_{2}-x_{1} }\]

OpenStudy (3psilon):

That will give you the Slope for the last two points

OpenStudy (anonymous):

Okay let me work it out tell me if its correct..

OpenStudy (3psilon):

So the slope would be \[\frac{ -8-7 }{ -2-3} = -15/-5 == \]

OpenStudy (3psilon):

Slope = 3

OpenStudy (3psilon):

And we know that parallel lines have the same slope

OpenStudy (anonymous):

I did it and got 3! thank you!

OpenStudy (3psilon):

Did you plug in the slope in y = mx + b form? and plug the first coordinates into that and solve for b?